OpenStudy (anonymous):

NEED HELP ASAP!!PLZ????!!!!!!! Variable y varies directly with x2, and y = 72 when x = 6. What is the value of y when x = 4? y = ? x^2

OpenStudy (anonymous):

o.O

OpenStudy (anonymous):

Im kinda in a hurry............

OpenStudy (anonymous):

Right, then let's make this quick. Can I assume you know what "varies directly" means?

OpenStudy (anonymous):

Not really :/

OpenStudy (anonymous):

Okay, two things vary directly if one of them if their QUOTIENT is always the same. So here, y varies directly with x^2. That means \[\Large \frac{x^2}{y}\] will ALWAYS be the same, no matter what values x and y take...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Now, we know it's always the same, we just need to know what IT is. You're given that if y is 72 then x is 6. Let's try plugging that in: \[\Large \frac{6^2}{72}\] What is this equal to?

OpenStudy (anonymous):

Common denominator is 12.... Simplify to : 1 whole? \[\frac{ 12 }{ 72 } \]

OpenStudy (anonymous):

Whoa... back up there :D \[\Large 6^2 \color{red}\ne 2\times 6\\\Large 6^2 = 6\times6\]

OpenStudy (anonymous):

Sorry my internet went out

OpenStudy (anonymous):

I did it wrong ???? @supreme_kurt

OpenStudy (anonymous):

I lost my connection too. I'm so sorry. :(